Have not gotten an answer yet...I sent it before but I really need help with this one :) Please!!! Someone help me out!
I have to use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
3z^2 + z - 1 = 0

Answer by stanbon(75887)

3z^2 + z - 1 = 0
b^2-4ac=1+12=13
Since the discriminant is positive
there are two Real Number roots.
Cheers,
Stan H.